given that in 1990, nominal GDP = 2063 and real GDP(base 1990) = 2063. in 1991, nominal GDP = 2113 and real GDP = 2088. for both years calculate:
a. nominal GDP growth
b. real GDP growth
c. GDP deflator
d. rate of inflation
1
Expert's answer
2014-11-05T12:07:48-0500
a. nominal GDP growth = nominal GDP (1991) / nominal GDP (1990) = 2113 / 2063 = 1.0242
b. real GDP growth = real GDP (1991) / real GDP (1990) = 2088 / 2063 = 1.0121
c. GDP deflator (1990) = nominal GDP (1990)/real GDP(1990) = 2063/2063 = 1
GDP deflator (1991) = nominal GDP (1991)/real GDP (1991) = 2113/2088=1.01197
d. rate of inflation (1990) = nominal GDP (1990)/ real GDP(1990) = 2063/2063 = 1
rate of inflation (1991) = nominal GDP (1991)/ real GDP(19910 = 2113/2088 = 1.0119